A firm’s balance sheet contains $3,200 of cash, $20,000 of fixed assets, $13,800 of accounts receivable, $7,200 of accounts payable, $5,000 of inventory, and $11,000 on a revolving line of credit. Assuming that list contains all of the company’s assets and liabilities:

11. How much are current assets? (Show calculation)

12. What is the company’s quick ratio? (Show calculation)

13. What is the company’s book value? (Show calculation)

PLEASE HELP THANK YOU IN ADVANCE
Answer #1

11

Current Assets = Cash + Cash Equivalents+ Inventory+ Accounts Receivables
Current Assets = +3200+5000+13800
Current Assets= 22000

12

Current Liabilities = Accounts Payable
Current Liabilities = 7200
Quick or acid test ratio = (current assets-inventory)/current liablities
Quick or acid test ratio = (22000-5000)/7200
Quick or acid test ratio = 2.36

13

Book value = Fixed asset+current assets-liabilities = 22000+20000-7200=34800

Note: Line of credit is not considered as it is an facility available to the company. It turns into a liability only when company avails this facility